# LLM-based Discriminative Reasoning for Knowledge Graph Question Answering

### v1

@@ Abstract
Large language models (LLMs) based on generative pre-trained Transformer have achieved remarkable performance on knowledge graph question-answering (KGQA) tasks. However, LLMs often produce ungrounded subgraph planning or reasoning results in KGQA due to the hallucinatory behavior brought by the generative paradigm, which may hinder the advancement of the LLM-based KGQA model. To deal with the issue, we propose a novel LLM-based Discriminative Reasoning (LDR) method to explicitly model the subgraph retrieval and answer inference process. By adopting discriminative strategies, the proposed LDR method not only enhances the capability of LLMs to retrieve question-related subgraphs but also alleviates the issue of ungrounded reasoning brought by the generative paradigm of LLMs. Experimental results show that the proposed approach outperforms multiple strong comparison methods, along with achieving state-of-the-art performance on two widely used WebQSP and CWQ benchmarks  [^1] Our code and data will be released..
@@ 1 Introduction
Large language models (LLMs) have shown remarkable reasoning capabilities in KGQA task Yu et al. (2022); Huang and Chang (2023); Wang et al. (2023b), especially the feasibility to prompt the LLMs to generate searching and reasoning results through the LLMs’ built-in knowledge. Typically, based on the given question, LLMs can be prompted to provide a plan for the question-related subgraph through one-time generation. After retrieving the subgraph, LLMs can directly generate the answers along with the reasoning steps using the subgraph as context. Utilizing internal knowledge or reasoning ability distilled from stronger models like GPT-4, the generative KGQA model can effectively conduct knowledge graph reasoning, along with achieving state-of-the-art performance on the KGQA tasks Mondorf and Plank (2024); LUO et al. (2024); Sun et al. (2024a).
@@
Figure 1: Example of previous generation-based methods and discriminative method proposed in this paper. The generation-based methods tend to generate unsupported or redundant subgraphs and reasoning results, while discriminative method address this issue by focusing only on the KG and the subgraph.
@@
Despite their success, the generative LLMs often produce ungrounded subgraph planning or reasoning results due to the hallucinatory behavior brought by the generation process Zhang et al. (2023); Sun et al. (2024b); Pan et al. (2024), which is opposite to the deterministic characteristic of knowledge reasoning process Garcez et al. (2015); Xiong et al. (2024). As shown in Figure 1, when searching question-related subgraphs, generation-based methods come with not existed path “1->5->3” due to hallucinatory planning (upper left), or retrieve redundant paths at one step “1->2; 1->3; 1->4” as a compensate to the generation uncertainty (upper middle). These issues lead to a noisy retrieved subgraph, which disperses the LLM’s attention during the reasoning process. When conducting answer inference on the retrieved subgraph, the generation-based methods may generate unreasonable reasoning steps “since r3" as inference chain or even entity “4” that do not exist in the subgraphs as answers (bottom left). As a result, the hallucinatory behavior of the generative LLMs hinders the performance of KGQA.
To address this issue, we propose a novel LLM-based Discriminative Reasoning (LDR) method to explicitly model the subgraph retrieval and answer inference process in the KGQA task. To this end, we reformulate the KGQA process into a discriminative reasoning framework, where KGQA is reformulated as three subtasks including subgraph searching, subgraph pruning, and answer inference. Thus, an LLM-based discriminative reasoning network is designed to model subgraph retrieval, pruning, and answer inference, thereby better achieving the KGQA. In summary, our main contributions are as follows:
  • We propose a discriminative reasoning framework to model the reasoning skills in KGQA. To the best of our knowledge, this is the first work to model different reasoning skills explicitly for enhancing the KGQA.
  • The proposed LDR organically combines the existing generative LLMs and our discriminative reasoning framework to advance the progress of KGQA tasks.
@@
  • Experimental results demonstrate that LDR achieved state-of-the-art performance on two widely used benchmarks.
2 Discriminative Reasoning Framework
@@
Figure 2: The proposed LDR of KGQA. Start from the question (bottom left) with a given starting entity "Coronation Street", LDR sequentially conducts subgraph retrieval, subgraph pruning, and answer inference. Then LDR automatically uses the reasoning results to prune $G_{k}$ and then retrieve the answers from it. In this figure, the node’s color in subgraph $G_{k}$ (middle center) represents its position in subgraph structure $S_{k}$ (top center).
@@
In this section, we propose a novel discriminative reasoning framework for KGQA which reformulates the KGQA task into three subtasks, including question-related subgraph searching, question-related subgraph pruning, and answer inference. Here we first formulate the KGQA task.
@@
Formulation of KGQA task. Given a question $Q$ and the knowledge graph entities $E$ contained in the question, the KGQA task asks the model to recall golden answers $A_{gold}$ as much as possible. The model has to retrieve question-related subgraphs from the knowledge graph $KG$ and infer the right answer based on the subgraphs.
@@
The knowledge graph $KG$ used in this paper is Freebase  [^2] The two benchmarks used in this work are constructed using Freebase Bollacker et al. (2008)., which consists of knowledge triplets represented as $t=(s,p,o)$ including the subject entity $s$, the object entity $o$, and the predicate $p$ that connects these two entities. Next, we will introduce the three subtasks in detail.
@@ 2.1 Question-related subgraph Searching
Formulation Given the knowledge graph, the question-related subgraph searching subtask aims to retrieve the question-related subgraph $G_{k}$ and thereby summarizes an abstract structure $S_{k}$ for each input question.
Specifically, $G_{k}$ is the subgraph comprising only the necessary knowledge to correctly answer the question $Q$, which can be represented by a set of triplets:
@@
The summarized abstract structure $S_{k}$ groups all the entities into “semantic nodes” based on their position in $G_{k}$:
@@ 2.2 Question-related Subgraph Pruning
Formulation Based on the abstract structure $S_{k}$, the question-related subgraph pruning task aims to map all the question-related constraint entities $C$ to nodes in $S_{k}$.
$C$ denotes the intersection of question mentioned entities $E_{question}$ and all entities in Freebase $E_{freebase}$:
@@ 2.3 Answer Inference
Formulation Given the question-related subgraph structure $S_{k}$, the answer inference sub-task aims to locate the position of the answer $A_{pos}$ in structure $S_{k}$.
The position of answers $A_{pos}$ corresponds to the position of $A_{gold}$ in $S_{k}$:
$A_{pos}=Grouped(A_{gold}),$ (5)
$A_{gold}\in Node(G_{k}),$
$A_{pos}\in Node(S_{k}).$
@@
3 LLM-based Discriminative Reasoning
Targeting three subtasks in the discriminative reasoning framework, we propose LLM-based discriminative reasoning (LDR) to explicitly complete the subtasks as illustrated in Figure 2.
3.1 Discriminative Searching
Previous methods rely on agent toolboxes add to the interaction overhead between the model and the knowledge graph and suffer from SPARQL syntax errors Cao et al. (2022); Gu et al. (2023); Jiang et al. (2024). Therefore, LDR opts for a discriminative searching approach which only observes and updates the subgraph structure $S_{k}$ during the process.
In each iteration, based on the retrieved subgraph structure $S_{k}$, the LLM selects one option from the option pool (as shown in Figure 2). Each option includes a starting node $s_{next}$ and a neighboring relation $p_{next}$, forming the next triple to search for:
$t_{next}=(s_{next},p_{next},o_{next}),$ (6)
$s_{next}\in Node(S_{k}).$
Since each node in the subgraph structure $S_{k}$ corresponds to a set of entities in $G_{k}$, conducting a search from a certain point in $S_{k}$ represents performing a search from each corresponding entity in $G_{k}$. A new node $o_{next}$ is added to $S_{k}$ along with its corresponding entities in $G_{k}$ retrieved from the knowledge graph. LDR maintains an option pool that includes all feasible triples for search. We can formulate each step of the discriminative searching strategy in LDR as:
@@ 3.1 Discriminative Searching
where $\mathbb{P}$ represents the probability distribution over options provided by the LLM using a constrained beam search algorithm based on output logits. After each decision made based on the option pool, LDR removes the selected option from the pool and expands the available neighboring directions introduced by the new entities. An additional option, ’None’, is always available to terminate the search process.
Retention of Node Information. LDR further enriches the information contained within the subgraph structure $S_{k}$ by labeling the semantic nodes with entity types. Entities in Freebase can be classified into one of the following types: entity, topic, date, and num (details are provided in Appendix A). LDR will recognize the type of retrieved entities, and use these to name newly added nodes in $S_{k}$.
3.2 Discriminative Pruning
After obtaining the question-related subgraph $G_{k}$ along with its structure $S_{k}$, LDR maps all constraints mentioned in the question onto $S_{k}$ to perform subgraph pruning. LDR determines the positions of constraints observing only the retrieved subgraph structure $S_{k}$, thereby isolating the influence of specific entity names in the task.
To apply a constraint $C_{n}$ to the subgraph, we need three elements: the constraint entity $c_{tar}$ mentioned in the questions, the target position $c_{pos}$ for applying the constraint, and the operator $c_{opt}$ to define the type of logical resolution used for applying the constraint: $C_{n}$ = ($c_{pos}$, $c_{opt}$, $c_{tar}$). We restrict the operator $c_{opt}$ to one of the seven types $\{=,<,\leq,>,\geq,\text{min},\text{max}\}$ and combine each constraint with all possible operators and positions to form the option pool.
Based on the question $Q$, structure $S_{k}$, LDR asks the LLM to iteratively select the constraints until the LLM selects ’None’ or there is no options left:
@@ 3.2 Discriminative Pruning
Once the pruning process is terminated by the LLM, LDR applies the constraint operations $C_{1},...,C_{n}$ to prune the subgraph $G_{k}$. The pruning process is conducted at the level of subtrees rooted from the starting node in $G_{k}$ (as shown in Figure 2), retaining only the subtrees that meet the constraints.
@@
UniKGQA (Jiang et al., 2022) 0.751 - 0.702 0.507 - 0.480
KD-CoT (Wang et al., 2023a) 0.686 - 0.525 0.557 - -
ToG w/ChatGPT (Sun et al., 2024a) 0.758 - - 0.589 - -
@@
LDR (Ours) 0.840 0.860 0.845 0.802 0.837 0.820
Table 1: The results of our method compared with previous approaches on WebQSP and CWQ. Asterisk * denotes the results we reproduced. Note that the Hits@1 result reported in the original RoG paper (WebQSP 0.857, CWQ 0.626) is not calculated in the right way, see the author’s response here.
3.3 Discriminative Answer inferring
When answering questions with a large number of answers, previous generative methods often fail to capture all the correct answers, even if the reasoning steps are successfully generated. To address this problem, LDR focuses on locating the positions of answers within the subgraph structure $S_{k}$ to simultaneously retrieve all possible answers. Based on $S_{k}$, LDR determines the answer position $A_{pos}$ using:
$A_{pos}=argmax(\mathbb{P}(n|S_{k},C,Q),n\in S_{k}).$ (9)
Here $\mathbb{P}$ is also given by the LLM based on $S_{k}$,$C$ and $Q$. Positions for applying constraints can not be chosen again. All entities in $G_{k}$ corresponding to the selected position $A_{pos}$ will be listed as the answer.
3.4 Constructing Training Data
To train the model, we utilize the SPARQL queries provided in the benchmarks. Here we briefly introduce the training data used for finetuning the LLMs in LDR.
Subgraph Searching. We make use of the SPARQL data available in existing benchmarks to form the training data. In WebQSP and CWQ, each question is associated with a SPARQL query. The direct execution of this query yields the answer to the open question (for details with SPARQL decomposition, see Appendix B). We obtain the correct subgraph structure required to solve each problem by decomposing the SPARQL statements. Unlike ROG (LUO et al., 2024), in finetuning process LDR always presents the model with the correct knowledge subgraph structure rather than the shortest path starts from the question entity and ends at the answer entities.
Subgraph pruning and answer inference. To finetune the LLMs to be capable of constraint determination and answer inference, we also construct constraint/answer locating samples from the SPARQL queries in WebQSP and CWQ. The input is a complete subgraph structure with all feasible options of constraints or answer positions, the golden output is the correct position of the constraint and the answer.
Data Statistics. Using the method described above, we construct our training set with 121,023 subgraph searching samples and 46,885 subgraph pruning and answer inference samples.
@@ 4.1 Datasets and Settings
Benchmarks. To evaluate the knowledge graph question-answering capability of the proposed method, we choose two widely used benchmarks, WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018) (for more details, see Appendix B).
Metrics. We choose commonly used metrics Hits@1 and F1 for the evaluation process following previous works LUO et al. (2024); Sun et al. (2024a); Jiang et al. (2024). For detailed definitions of metrics, please refer to Appendix C.
Baselines. We compare LDR with previous SOTA generation-based KGQA strategies including RoG with LLM planning and chain-of-thought reasoning strategy (LUO et al., 2024), ToG with interactive reasoning strategy (Sun et al., 2024a) and KGAgent with tool-based agent reasoning strategy (Jiang et al., 2024). We also list previous representative methods and zero-shot performances of three widely used LLMs for comparison.
Base Model. To ensure fairness in comparison, we choose Llama2-7b (Touvron et al., 2023) as the base model of LDR following RoG (LUO et al., 2024) and KGAgent (Jiang et al., 2024).

### v2

@@ Abstract
Large language models (LLMs) based on generative pre-trained Transformer have achieved remarkable performance on knowledge graph question-answering (KGQA) tasks. However, LLMs often produce ungrounded subgraph planning or reasoning results in KGQA due to the hallucinatory behavior brought by the generative paradigm. To tackle this issue, we propose READS to reformulate the KGQA process into discriminative subtasks, which simplifies the search space for each subtasks. Based on the subtasks, we design a new corresponding discriminative inference strategy to conduct the reasoning for KGQA, thereby alleviating hallucination and ungrounded reasoning issues in LLMs. Experimental results show that the proposed approach outperforms multiple strong comparison methods, along with achieving state-of-the-art performance on widely used benchmarks WebQSP and CWQ.  [^1] Our code and data will be released upon acceptance.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable reasoning capabilities in KGQA task Yu et al. (2022); Huang and Chang (2023); Zhu et al. (2024), especially the feasibility to prompt the LLMs to generate searching and reasoning results through the LLMs’ built-in knowledge. Typically, based on the given question, LLMs can be prompted to provide a plan for the question-related subgraph through one-time generation. After retrieving the subgraph, LLMs can directly generate the answers along with the reasoning steps using the subgraph as context. Utilizing internal knowledge or reasoning ability distilled from stronger models like GPT-4, the generative KGQA model can effectively conduct knowledge graph reasoning, along with achieving state-of-the-art performance on the KGQA tasks Mondorf and Plank (2024); LUO et al. (2024); Sun et al. (2024a).
@@
Figure 1: The generation-based methods tend to generate unsupported or redundant subgraphs and reasoning results (left), while the proposed method address the issue by establishing proper searching space for each of the KGQA subtasks (right).
@@
Despite their success, the generative reasoning methods often produce ungrounded planning or reasoning results due to the hallucinatory behavior Zhang et al. (2023); Sun et al. (2024b); Pan et al. (2024), which is opposite to the deterministic characteristic of knowledge reasoning process Garcez et al. (2015); Xiong et al. (2024). As shown in Figure 1, when searching question-related subgraphs, generation-based methods come with not existed path “1->5->3” due to hallucinatory planning , or retrieve redundant paths at one step “1->2; 1->3; 1->4” as a compensate to the generation uncertainty (upper left). When conducting answer inference on the retrieved subgraph, the generation-based methods may generate unreasonable step “since r3" as inference chain or even entity “4” that do not exist in the subgraphs as answers (bottom left). The hallucinatory behavior of the generative LLMs hinders the advancement of KGQA.
To address the issue, we propose LLM-Based Reasoning With Discriminative Subtasks (READS) to strengthen the LLM-based knowledge reasoning process. READS decomposes the KGQA process into three discriminative subtasks: graph searching, graph pruning, and answer inference. The decomposition aims to explicitly simulate the capabilities of searching for question-related knowledge, identifying semantic constraints, and inferring the answer position on the subgraph, respectively. Meanwhile, READS simplifies search space from the knowledge graph without toolboxes, along with designed discriminative inference strategy to conduct the reasoning of KBQA effectively. In summary, our main contributions are as follows:
  • We introduce READS, an novel reasoning framework that explicitly models KGQA reasoning skills by deconstructing the KGQA process into three discriminative subtasks.
  • An effective corresponding discriminative inference strategy is designed to conduct the reasoning of KGQA for READS, thereby significantly alleviating hallucination and ungrounded reasoning issues.
@@
  • Experimental results demonstrate that READS achieved state-of-the-art performance on two widely used benchmarks.
2 Related Works
Generative Approaches. The challenge of KGQA task lies in how to conduct precise reasoning on the knowledge graphs Miller et al. (2016); Bai et al. (2022); Zhang et al. (2024), early works tried to teach models to construct database queries for knowledge graphs, allowing them to directly retrieve answers from the graph Gu and Su (2022); Ye et al. (2022). With the advent of LLM’s long-horizon planning and reasoning capability Zhong et al. (2024); Wang et al. (2024), the focus of KGQA research shifts toward leveraging the reasoning capabilities of a single LLM for knowledge inference Jiang et al. (2022). One straightforward way is to directly schedule the question-related sub-graph using the LLM’s knowledge Hong et al. (2023); Wang et al. (2023). Typical approach like RoG employs chain-like subgraph planning and distills GPT-4’s Chain-of-Thought reasoning capability to achieve reliable reasoning processes over knowledge graphs, achieving state-of-the-art performance (LUO et al., 2024). Despite their success, one concern is that those methods often provides incorrect and ungrounded reasoning results due to the hallucinatory generation process.
Interactive and Discriminative Approaches. An alternative approach is to design effective tools and generation strategies to retrieve environment information from the knowledge graph to enhance the step-by-step reasoning process, as seen in approaches such as ToG, KGAgent and GoG Sun et al. (2024a); Jiang et al. (2024); Xu et al. (2024). However, these interactive generation approaches cannot avoid the influence of hallucinations. Even when provided with environmental information or recalled subgraphs, the model may still arrive at incorrect reasoning results. As claimed in PANGU, using discriminative strategy can effectively mitigates the hallucination problem Gu et al. (2023). Despite PANGU’s success, integrating tools such as search and answer retrieval within the same search space may also lead the LLM to make erroneous decisions.
In this paper, we propose to reformulate KGQA into three subtasks to explicitly model the KGQA skills and design discriminative strategies to effectively enhance LLM’s reasoning capability on the knowledge graph.
3 READS Framework
@@
Figure 2: The proposed READS for KGQA. Start from the question (bottom left) with a given starting entity "Coronation Street", READS sequentially conducts subgraph retrieval, subgraph pruning, and answer inference. Then READS automatically uses the reasoning results to prune $G_{k}$ and then retrieve the answers from it. In this figure, the node’s color in subgraph $G_{k}$ (middle center) represents its position in subgraph structure $S_{k}$ (top center).
@@
In this section, we propose a novel framework that reformulates the KGQA task into three discriminative subtasks, including question-related subgraph searching, question-related subgraph pruning, and answer inference.
@@
Formulation of KGQA task. Given a question $Q$ and the knowledge graph entities $E$ contained in the question, the KGQA task asks the model to recall golden answers $A_{gold}$ as much as possible. The model has to retrieve question-related subgraphs from the knowledge graph and infer the right answer based on the subgraphs.
@@
The knowledge graph $KG$ used in this paper is Freebase [^2] The two benchmarks used in this work are constructed using Freebase Bollacker et al. (2008)., which consists of knowledge triplets represented as $t=(s,p,o)$ including the subject entity $s$, the object entity $o$, and the predicate $p$ that connects these two entities.
@@ 3.1 Question-related subgraph Searching
Given the knowledge graph, the question-related subgraph searching subtask aims to retrieve the question-related subgraph $G_{k}$ and thereby summarizes an abstract structure $S_{k}$ for each input question. Specifically, $G_{k}$ is the subgraph comprising only the necessary knowledge to correctly answer the question $Q$, which can be represented by a set of triplets:
@@
Following UniKGQA Jiang et al. (2022), we use “semantic nodes” to represent a group of entities sharing same structural position in the knowledge graph. The summarized abstract structure $S_{k}$ groups all the entities into “semantic nodes” based on their position in $G_{k}$:
@@ 3.2 Question-related Subgraph Pruning
Based on the abstract structure $S_{k}$, the question-related subgraph pruning task aims to map all the question-related constraint entities $C$ to nodes in $S_{k}$. $C$ denotes the intersection of question mentioned entities $E_{question}$ and all entities in Freebase $E_{freebase}$:
@@ 3.3 Answer Inference
Given the question-related subgraph structure $S_{k}$, the answer inference subtask aims to locate the position of the answer $A_{pos}$ which corresponds to the position of $A_{gold}$ in $S_{k}$:
$A_{pos}=\text{Grouped}(A_{gold}),$ (5)
$A_{gold}\in\text{Node}(G_{k}),$
$A_{pos}\in\text{Node}(S_{k}).$
@@
4 READS Discriminative Reasoning
After we propose the framework of READS, we are able to design efficient reasoning strategy to facilitate graph retrieval, graph pruning and answer inference. Based on the subtasks, we are able to explicitly model the KGQA process as illustrated in Figure 2. We design discriminative strategies to achieve the subtasks and construct training data to augment the LLM-based reasoning process.
4.1 Searching Strategy
Compared to previous methods that rely on agent toolboxes, READS opts for a discriminative searching approach which only observes and updates the subgraph structure $S_{k}$.
In each iteration, based on the retrieved subgraph structure $S_{k}$, the LLM selects one option from the option pool (as shown in Figure 2). Each option includes a starting node $s_{next}$ in $S_{k}$ and a neighboring relation $p_{next}$, forming the next triple $t_{next}=(s_{next},p_{next},o_{next})$. A new node $o_{next}$ is added to $S_{k}$ along with its corresponding entities in $G_{k}$ retrieved from the knowledge graph. READS maintains an option pool that includes all feasible triples for search. We can formulate each step of the discriminative searching strategy in READS as:
@@ 4.1 Searching Strategy
where $\mathbb{P}$ represents the probability distribution over options provided by the LLM using a constrained beam search algorithm based on output logits. An additional option, ‘None’, is always available to terminate the search process.
Retention of Node Information. READS further enriches the information contained within the subgraph structure $S_{k}$ by labeling the semantic nodes with entity types. Entities in Freebase can be classified into one of the following types: entity, topic, date, and num (details are provided in Appendix A). READS will recognize the type of retrieved entities, and add new nodes in $S_{k}$.
4.2 Pruning Strategy
After obtaining the question-related subgraph $G_{k}$ along with its structure $S_{k}$, READS maps all constraints mentioned in the question onto $S_{k}$ to perform subgraph pruning. A triplet $C_{n}$ = ($c_{pos}$, $c_{opt}$, $c_{tar}$) has to be chosen from the option pool, where $c_{tar}$ is the constraint entity mentioned in the question, $c_{pos}$ is the target position for applying the constraint, and $c_{opt}$ is the operator to define the type of logical resolution used for applying the constraint. READS restrict the operator $c_{opt}$ to one of the seven types $\{=,<,\leq,>,\geq,\text{min},\text{max}\}$ and combine each constraint with all possible operators and positions to form the option pool.
Based on the question $Q$ and $S_{k}$, READS asks the LLM to iteratively select the constraints until the LLM selects ‘None’ or there are no options left:
@@ 4.2 Pruning Strategy
The pruning process is conducted at the level of subtrees rooted from the starting node in $G_{k}$ (as shown in Figure 2), retaining only the subtrees that meet the constraints.
@@
GPT-4-turbo zero-shot Achiam et al. (2023)* 0.632 - - 0.483 - -
Llama2-7b SPARQL Generation* 0.747 - - 0.656 - -
@@
UniKGQA (Jiang et al., 2022) 0.772 - 0.722 0.512 - 0.490
PANGU (Gu et al., 2023) 0.796 - - 0.622 - -
KD-CoT (Wang et al., 2023) 0.686 - 0.525 0.557 - -
@@
READS (Ours) 0.840 0.860 0.845 0.802 0.837 0.820
Table 1: The results of our method compared with previous approaches on WebQSP and CWQ. Asterisk (*) denotes the results we reproduced. Note that the Hits@1 result reported in the original RoG paper (WebQSP 0.857, CWQ 0.626) is not calculated in the right way, see the author’s response here.
4.3 Answering Strategy
When answering questions with a large number of answers, previous generative methods often fail to capture all the correct answers, even if the reasoning steps are successfully generated. To address this problem, READS focuses on locating the positions of answers within the subgraph structure $S_{k}$ to simultaneously retrieve all possible answers. Based on $S_{k}$, READS determines the answer position $A_{pos}$ using:
$A_{pos}=argmax(\mathbb{P}(n|S_{k},C,Q),n\in S_{k}),$ (8)
where $\mathbb{P}$ is also given by the LLM based on $S_{k}$,$C$, and $Q$. Positions for applying constraints can not be chosen again. All entities in $G_{k}$ corresponding to the position $A_{pos}$ will be listed as the answer.
@@ 5.1 Datasets and Settings
Data preprocessing. Based on the proposed subtasks, we construct training data based on the original training sets. We get 121,023 subgraph searching samples and 46,885 subgraph pruning and answer inference samples. For more details of our data preprocessing and training data construction method, please refer to Appendix B.
Benchmarks. To evaluate the knowledge graph question-answering capability of the proposed method, we choose two widely used benchmarks, WebQSP (Yih et al., 2016) and CWQ (Talmor and Berant, 2018). These two benchmarks are constructed based on Freebase knowledge graph.
Metrics. We choose commonly used metrics Hits@1 and F1 for the evaluation process following previous works LUO et al. (2024); Sun et al. (2024a). For detailed definition and implementation of the metrics, please refer to Appendix C.
Baselines We use previous reproducible SOTA generation-based KGQA method RoG as our baseline. RoG make full use of LLM planning and chain-of-thought reasoning capability to achieve remarkable KGQA performance LUO et al. (2024). We also listed typical methods like ToG and KGAgent with interactive reasoning strategy (Sun et al., 2024a; Jiang et al., 2024), PANGU with single-task discriminative strategy Gu et al. (2023). The zero-shot performance of widely used LLMs is listed for comparison. We also finetuned llama2-7b to directly generate SPARQL queries for each of the question, and then execute those queries on Freebase to get the answer.
Base Model. We choose Llama2-7b as the base model of READS following RoG. For implementation with GPT-4, see Section 5.7.